OpenStudy (anonymous):

Calculus: Find the derivative. Where does this function have horizontal tangents? y=x^2/(x-2)

OpenStudy (anonymous):

\[y=\frac{ x ^{2} }{ x-2 }\]

hartnn (hartnn):

do u know how to get the slope of tangent to a curve ?

OpenStudy (anonymous):

make y 0 rite? :)

OpenStudy (anonymous):

and then find the derivative of this equation?

hartnn (hartnn):

to get slope of tangent to the curve, you differentiate y to get y' now horizontal tangent implies slope = 0 , so then put y'=0 not y=0

hartnn (hartnn):

so the order is 1st find the derivative, then put y'=0

OpenStudy (anonymous):

\[\frac{ x ^{2} -4}{ (x-2)^{2} }\]

OpenStudy (anonymous):

thats what i got when use the quotient rule to find the derivative

hartnn (hartnn):

wouldn't that be 4x ? instead of 4 ?

OpenStudy (anonymous):

yeah sorry my cp froze on me xD

hartnn (hartnn):

so now equate that to 0

OpenStudy (anonymous):

\[\frac{ x ^{2}-4x }{ (x-2)^{2} }\]

OpenStudy (anonymous):

ok then what :)

hartnn (hartnn):

that is your y' equate that to 0, what u get ?

OpenStudy (anonymous):

so.. \[\frac{ x ^{2}-4x }{ (x-2)^{2} }=0\]

hartnn (hartnn):

yes, which gives you ?

OpenStudy (anonymous):

idk. :3 lol

hartnn (hartnn):

when a fraction = 0, its numerator = 0

OpenStudy (anonymous):

yeap so then what do i do with this equation

hartnn (hartnn):

so u get \(x^2-4x=0\) so which are 2 values of x, that make x^2-4x=0 ?

OpenStudy (anonymous):

0 and 4? :D

hartnn (hartnn):

that is correct

OpenStudy (anonymous):

ok so my x values are 0,4 how do i find my y values

hartnn (hartnn):

put it in y=.....

OpenStudy (anonymous):

do i plug in my x values into the denominator to get my y values?

hartnn (hartnn):

put x= 0 here \(\large y=\frac{ x ^{2} }{ x-2 }\) what u get ?

OpenStudy (anonymous):

oh so in the original equation?

hartnn (hartnn):

yes :)

hartnn (hartnn):

so which two points did u get ?